A sample of 10 is taken from a days output of a machine that produces parts of which 5% are defective. A full inspection of a days production is conducted whenever the sample of 10 gives 2 or more defective parts. Find the probability of a full inspection? What assumptions do you need to make?
Answer #1

here this is is binomial distribution for which parameter n=10 and p=0.05

we are assuming that sample is random ; and probability of one product being defective is independent of other product.

P(full inspection)=P(2 or more defective)=P(X>=2)=1-P(X<=1)=1-(P(X=0)+P(X=1))

=1-10C0(0.05)0(0.95)10-10C1(0.05)1(0.95)9 =0.0861